# Knots and variance ordering of sequential Monte Carlo algorithms

### v1

@@ Appendix A Supporting results
as $M_{p}^{G_{p}}=R_{p}^{K_{p}(G_{p})}K_{P}^{G_{p}}$ by Proposition A.3 and since $R_{p}K_{p}=M_{p}$. Finally, let $\mathcal{K}^{\prime}=(R_{0:n-1}^{\prime},K_{0:n-1}^{\prime})$ and $\mathcal{M}_{2}=\mathcal{K}^{\prime}\ast\mathcal{M}^{\diamond}$ where $R_{0}^{\prime}=\delta_{0}$ and $K_{0}^{\prime}=\mathrm{Id}$, $R_{p}^{\prime}=R_{p-1}^{K_{p-1}(G_{p-1})}$ and $K_{p}^{\prime}=K_{p-1}^{G_{p-1}}$ for $p\in[1\,{:}\,n-1]$. Therefore from Proposition 3.9, $\mathcal{M}_{2}=(M_{2,0:n},G_{2,0:n})$ satisfies $M_{2,p}=M_{1,p}$ and $G_{2,p}=G_{1,p}$ for $p\in[0\,{:}\,n]$ and hence $\mathcal{M}_{2}=\mathcal{M}_{1}$ completing the proof. ∎
Proposition A.7 (Repeated adapted knotset equivalence).
@@
The proof follows from repeated applications of Proposition A.6. ∎
Proposition A.8 (Terminal knot kernel equivalence).
@@
Then, by Proposition A.1, $K(H)\cdot K^{H}P_{2}(\phi^{-1}\cdot\varphi)=K\{H\cdot P_{2}(\phi^{-1}\cdot\varphi)\}$ and using compatibility condition (ii) we have

### v2

@@ Appendix B Supporting results
as $M_{p}^{G_{p}}=R_{p}^{K_{p}(G_{p})}K_{P}^{G_{p}}$ by Proposition B.3 and since $R_{p}K_{p}=M_{p}$. Finally, let $\mathcal{K}^{\prime}=(R_{0:n-1}^{\prime},K_{0:n-1}^{\prime})$ and $\mathcal{M}_{2}=\mathcal{K}^{\prime}\ast\mathcal{M}^{\diamond}$ where $R_{0}^{\prime}=\delta_{0}$ and $K_{0}^{\prime}=\mathrm{Id}$, $R_{p}^{\prime}=R_{p-1}^{K_{p-1}(G_{p-1})}$ and $K_{p}^{\prime}=K_{p-1}^{G_{p-1}}$ for $p\in[1\,{:}\,n-1]$. Therefore from Proposition 3.9, $\mathcal{M}_{2}=(M_{2,0:n},G_{2,0:n})$ satisfies $M_{2,p}=M_{1,p}$ and $G_{2,p}=G_{1,p}$ for $p\in[0\,{:}\,n]$ and hence $\mathcal{M}_{2}=\mathcal{M}_{1}$ completing the proof. ∎
Proposition B.7 (Repeated adapted knotset equivalence).
@@
The proof follows from repeated applications of Proposition B.6. ∎
Proposition B.8 (Terminal knot kernel equivalence).
@@
$=$ $R\{K[H\cdot P_{2}(\phi^{-1}\cdot\varphi)]\},$
where the last equality follows from Proposition B.1. Finally, using Definition 6.3 compatibility condition (ii) we have